OpenStudy (anonymous):

A 10-foot ladder leans against a wall with its foot braced 3 feet from wall's base. How far up the wall does the ladder reach? Pls help..:(

OpenStudy (anonymous):

use Pythagoras's theorem (10)^2=(3)^2+(h)^2 100-9=h^2 h=9.54

OpenStudy (anonymous):

Thank you so much for ur help math is not my thing I truly appreciated:)

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

U think u could help me on my other questions if its ok with u..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Multiply or divide. 3n^2-n/n^2-1divided by n^2/n+1

OpenStudy (anonymous):

(3n^2-n/n^2-1)/(n^2/n+1) (3n^2-n/n^2-1)x(n+1/n^2) (n(3n-1)/(n+1)(n-1))x(n+1/n^2) n gets cancelled with n^2 n+1 gets cancelled with n+1 remaining n is multiplied with the denominator (3n-1)/(n(n-1))

OpenStudy (anonymous):

is it ok?

OpenStudy (anonymous):

Yea it's good what abt this one? Find the value of each expression.9p3 and the other one is 10c4 thnx...

OpenStudy (anonymous):

9P3 = Permutations any three from 9 = 9! / (9 -3)! = 9! / 6! = 9 * 8 * 7 = 504 10C4 = Combinations any 4 from 10 = 10! / 4! (10 - 4)! = 10! / 4!6! = (10*9*8*7) / (4*3*2*1) = 5040 / 24 = 210 don't forget to use the best response button :)

OpenStudy (anonymous):

There are 15 members of the show choir. In how many ways can you arrange 4 members in the front row?

OpenStudy (anonymous):

What is the response button?

OpenStudy (anonymous):

15P4 =32760

OpenStudy (anonymous):

Can u show how u get the answer pls thnx

OpenStudy (anonymous):

15!/(15-4)! 15!/11! (11!x12x13x14x15)/11! 11 factorial cancels out 12x13x14x15=32760

OpenStudy (anonymous):

Multiply or divide.5x+10/x+2•2x/4x-10

OpenStudy (anonymous):

I just have 3 more problem if u can help me I would truly appreciate ur help thank you for all the help...

OpenStudy (jhannybean):

Are you also attempting them yourself? :)

OpenStudy (anonymous):

10x^2+20x/(4x^2-10x+8x-20) 10x^2+20x/(4x^2-2x-20) 5x^2+10x/(2x^2-x-10)

OpenStudy (anonymous):

Do I have to cancel anything to get the final answer?

OpenStudy (anonymous):

no only take 2 common from numerator and denominator which then cancel out

OpenStudy (anonymous):

Simplify each radical expression. Sqrt 75+Sqrt 3...next one is Sqrt 7(Sqrt 14+Sqrt 3).

OpenStudy (anonymous):

not this time post your solutions and I will correct them

OpenStudy (anonymous):

can u do it or not??

OpenStudy (anonymous):

No not really can u show me how to do it pls

OpenStudy (anonymous):

sqrt(75)+sqrt(3) =sqrt(25x3)+sqrt(3) 5sqrt(3)+sqrt(3) 6sqrt(3) get it?

OpenStudy (anonymous):

sqrt(7)(sqrt(7x2)+sqrt(3)) 7sqrt(2)+sqrt(3)

OpenStudy (anonymous):

Thnx very much...

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

oh sorry I missed something the answer is 7sqrt(3)+sqrt(7x3) =7sqrt(3)+sqrt(21)

OpenStudy (anonymous):

Thank you so much again for help I wish some times I'm good in math like u guys out there...